OpenStudy (anonymous):

What are the equations of the other planes of a 4X4X4 cube, given that the equation of one of the planes is 3x + 2y - z =4? I haven't worked with vectors in 5 years and my sister needs help with this. I have an idea how to do it, but I'm not sure my equations are right.

OpenStudy (amistre64):

cubes have parallel sides and perpendicular sides right?

OpenStudy (amistre64):

one of the sides will have to be a distance of 4 from this plane, you know what a normal is?

OpenStudy (anonymous):

Yes, but can I just change the 4 to 8 or 0 to get a parallel line with distance 4 away?

OpenStudy (amistre64):

i do not believe so, you have to measure along the normal vector; not just randomly add stuff. it might work but i cant guarentee it :)

OpenStudy (anonymous):

It doesn't work by changing the 4, I wolfram-ed it.

OpenStudy (anonymous):

How would you do find the parallel plane?

OpenStudy (amistre64):

take a point on the plane, and add 4 unit normals to it to find a point in t parallel plane that is 4 away

OpenStudy (amistre64):



OpenStudy (amistre64):

3x + 2y - z =4, a point on the plane; let x=0, y=0, z has to be -4

OpenStudy (amistre64):

or, let x=0, z=0, y has to 2

OpenStudy (amistre64):

how do we know the normal, and how do we unit it?

OpenStudy (anonymous):

normal is [3,2,-1] unit is [3,2,-1]/sqrt(14)

OpenStudy (amistre64):

good now we just add vectors (0,2,0) + (3,2,-1)/sqrt(14)

OpenStudy (amistre64):

forgot to times 4 lol

OpenStudy (amistre64):

(0,2,0) + 4(3,2,-1)/sqrt(14) (0,2,0) + (12,8,-4)/sqrt(14)

OpenStudy (amistre64):

1/sqrt(14) = sqrt(14)/14 ifn you want to simplify things ... dunno how simple they get tho

OpenStudy (anonymous):

That makes sense

OpenStudy (amistre64):

(0,2,0) + sqrt(14)(12,8,-4)/14 (0,2,0) + sqrt(14)(6,4,-2)/7 add like parts point on parallel plane is: 6sqrt(14)/7 , 2 + 4sqrt(14)/7 , -2sqrt(14)/7 a b c now, the equation of the new plane just differs by a constant, 3x +2y -z = K, such that K=3a +2b -c

OpenStudy (anonymous):

What about adding 4 unit vectors to the original equation 3(x-12/sqrt(14)) + 2(y-8/sqrt(14)) - (z+4/sqrt(14))=4

OpenStudy (amistre64):

im not sure that would work, we can test it out but id work the confident way first to make sure :)

OpenStudy (anonymous):

Ok I'll do that

OpenStudy (amistre64):

3(6sqrt(14)/7)+2(2 + 4sqrt(14)/7)-(-2sqrt(14)/7) = 4(1+sqrt(14)) which is about 19 if we round but we dont want to round do we

OpenStudy (amistre64):

now comes a question ive been wondering since we started .... does the cube give us any placement of it? if we only know 1 side of it; but not anything else, then were are the corners sopse to be? do we just randomly make one thats 4x4x4?

OpenStudy (amistre64):

good news is, my approach and your shifting approach work the same results :)

OpenStudy (anonymous):

I think we just randomly work out any cube with 4X4X4

OpenStudy (amistre64):

then we should decide on a vertex point to establish the other planes with

OpenStudy (amistre64):

3x +2y -z = 4 we have 3 to play with that are the intercepts 0,0,-4 0,2,0 4/3,0,0

OpenStudy (anonymous):

0,2,0 seems good

OpenStudy (amistre64):

hmm, then it must be a trap lol

OpenStudy (amistre64):

lets ride along the vertex from 0,2,0 towards 0,0,-4 since we know they are in the plane and we can make a direction vector from them

OpenStudy (anonymous):

It's not. It's actually weird. My sis came to me and showed me a picture that was on her class's whiteboard - the teacher wrote "on tuesday, you'll be given the equation of a plane and asked to create other planes to build a geometric shape in 3d. you will have to show your planes work. you may have dimensions specified"

OpenStudy (anonymous):

So my sister and I made up a very general question and I'm trying to work from that to give her an idea of what to do

OpenStudy (anonymous):

Ok, makes sense

OpenStudy (amistre64):

your doing fine. well cover what we did in the end in a general after we have established a good path to take

OpenStudy (amistre64):

0,0,-4 - 0,2,0 = 0,-2,-4, right? this should be our direction vector from 0,2,0

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

unit it and times it by 4

OpenStudy (amistre64):

and cross it with the normal to get the vector in the other direction ...

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

n x v = p such that p is perp to n and v these vectors will be the 'normals' to their respective sides

OpenStudy (amistre64):

and since they are 4 from 0,2,0 the will also provide points for anchoring

OpenStudy (amistre64):